OpenStudy (anonymous):

Simplify square root parenthesis 1 minus cosine theta parenthesis times parenthesis 1 plus cosine theta parenthesis. ±sin Θ ±cos Θ ±tan Θ square root cosine theta

OpenStudy (e.mccormick):

\(\sqrt{-\cos\theta }\cdot(1+\cos\theta)\) ? btw, I used this to type that: ``` \(\sqrt{-\cos\theta }\cdot(1+\cos\theta)\) ```

OpenStudy (e.mccormick):

Ooops. Missed a 1 there.

OpenStudy (e.mccormick):

\(\sqrt{1-\cos\theta }\cdot(1+\cos\theta)\)

OpenStudy (anonymous):

correct

OpenStudy (e.mccormick):

Hmmm. Not sure how much that would simplify.

OpenStudy (anonymous):

yea its trig it sucks :c

OpenStudy (e.mccormick):

Yah, I don't see a half angle that quite mathes it.

OpenStudy (e.mccormick):

You could do the distribution, but I do not know if that will really do much.

OpenStudy (dan815):

mccccormickk where have u beeeen

OpenStudy (e.mccormick):

Hiding. You have any good ideas on this one?

OpenStudy (dan815):